Basic bond valuation Complex Systems has an outstanding issue of ​$1,000​-par-value bonds with a 15​% coupon interest rate. The issue pays interest annually and has 16 years remaining to its maturity date.

a.  If bonds of similar risk are currently earning a rate of return of 11​%, how much should the Complex Systems bond sell for​ today?  

b.  Describe the two possible reasons why the rate on​ similar-risk bonds is below the coupon interest rate on the Complex Systems bond.

c.  If the required return were at 15​% instead of 11​%, what would the current value of Complex​ Systems' bond​ be? Contrast this finding with your findings in part a and discuss.
Answer #1

a. YTM =11%
Number of years =16
Par Value =1000
Coupon =15%*1000 =150
Price of Bond Today =PV of Coupons +PV of Par Value =150*((1-(1+11%)^-16)/11%+1000/(1+11%)^16 =1295.17

b. Two reasons why interest rate is below coupon rate
1. The bond might be a company with higher credit rating . Hence YTM is lower than coupon rate.
2. The leverage or debt ratio of the company might be very low. Lower the leverage, lower the risk and lower the YTM.

c. YTM =15%
Number of years =16
Par Value =1000
Coupon =15%*1000 =150
Price of Bond Today =PV of Coupons +PV of Par Value =150*((1-(1+15%)^-16)/15%+1000/(1+15%)^16 =1000.00

Higher the YTM lower the price and lower the YTM higher the price